40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Nov 20 15:11:09 2011, kiwamu, Update, IOO, RFAM monitoring test 
    Reply  Mon Nov 21 12:48:00 2011, Jenne, Update, IOO, RFAM monitoring test 
    Reply  Mon Nov 21 14:15:25 2011, Jenne, Update, IOO, RFAM monitoring test 
    Reply  Mon Nov 21 14:35:28 2011, kiwamu, Update, IOO, RFAM monitoring test REFL_RFAM.png
       Reply  Mon Nov 21 17:48:36 2011, Koji, Update, IOO, RFAM monitoring test 
Message ID: 5968     Entry time: Mon Nov 21 14:35:28 2011     In reply to: 5964     Reply to this: 5972
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: RFAM monitoring test 

REFL_RFAM.png

 This is a trend for a day long showing the REFL11/55 demod signals, REFLDC (corresponding to the MC transmitted power) and the PSL booth temperatire.

There are sudden jumps in the REFL55_I and REFL11_Q signals around 5:00 AM this morning, also at the same time the temperature suddenly went up.

But the quality of the signal turned out to be not so good because the fluctuation is still within 1 bit of the ADCs,

we have to try it again with a bigger gain in the analog whitening circuit.

Quote:

An RFAM measurement is ongoing

ELOG V3.1.3-